Mr. Warren drove from Twonsville to Villaville and back again at the speeds shown. Hit total driving time was 12 hours. How far apart are the two towns?
-----> 70mph
<----- 50mph

Answer by Alan3354(69443)

Mr. Warren drove from Twonsville to Villaville and back again at the speeds shown. Hit total driving time was 12 hours. How far apart are the two towns?
-----> 70mph
<----- 50mph
--------------
Avg speed for the round trip = 2*70*50/(70+50)
= 700/12 = 175/3 mi/hr
----
TR distance = r*t = (175/3)*12
= 700 miles
--> 350 miles each way